# Interpretable Company Similarity with Sparse Autoencoders

### v1

@@ Abstract
Determining company similarity is a vital task in finance, underpinning hedging, risk management, portfolio diversification, and more. Practitioners often rely on sector and industry classifications to gauge similarity, such as SIC-codes and GICS-codes – the former being used by the U.S. Securities and Exchange Commission (SEC), and the latter widely used by the investment community. Clustering embeddings of company descriptions has been proposed as a potential technique for determining company similarity, but the lack of interpretability in token embeddings poses a significant barrier to adoption in high-stakes contexts. Sparse Autoencoders (SAEs) have shown promise in enhancing the interpretability of Large Language Models (LLMs) by decomposing LLM activations into interpretable features. In this paper, we explore the use of SAE features in measuring company similarity and benchmark them against (1) SIC codes and (2) Major Group codes. We conclude that SAE features can reproduce and even surpass sector classifications in quantifying fundamental characteristics of companies, evaluated by the correlation of monthly returns, a proxy for similarity, and PnL from cointegration.
@@
Marco Molinari*, 1and Vladimir Tregubiak*, 1and Victor Shao*, 1 Abhimanyu Pandey1  and Mateusz Mikolajczak1  and Sebastião Kuznetsov Ryder Torres Pereira1 1 LSE.AI, London School of Economics and Political Science https://lseai.org Correspondence: m.molinari1@lse.ac.uk *Equal Contribution
@@ 1 Introduction
Traditional modes of comparison between companies are based on relative returns, and, in a discrete fashion, the sector sub-category. This is based on classification frameworks such as the Global Industry Classification System (GICS), which splits companies into 11 sectors and additional sub-categories for a total of 163 sub-industries (MSCI, 2020). Relying on the spread between returns can be effective but is not full-proof, since volatile periods in the market, changes in economic conditions, and temporal considerations can break spreads (Loretan and English, 2000).
Additionally, GICS classification is limited as the restricted granularity of a discrete classification system makes it difficult to obtain a dynamic interpretation of a company’s nature (Winton, 2018). This system falls short of taking into account the duality of certain companies, particularly in emerging industries (Winton, 2018; Anderson, 2018).
This is particularly important for Pairs Trading, a market-neutral strategy structured around the mean-reverting assumption of spreads between returns (Ehrman, 2012). Employing a pairs trading strategy with fundamentally similar companies whose returns are co-integrated could lessen the risk of deviation from historical trends (Bharadwaj, 2014).
Clustering embeddings of company descriptions has been proposed as a measure of similarity (Vamvourellis et al., 2023), but token embeddings are not interpretable, and this uncertaninty is undesirable in the financial sector.
SAEs have the potential to provide a more efficient measure of company similarity by allowing large amounts of financially and strategically relevant data to be decomposed into interpretable and meaningful features (Chen et al., 2020). This paper posits that employing SAEs as a method for company comparison is also practically efficient, as the method heavily alleviates the need for human input and analysis which is typical of traditional financial methods.
@@ 1.1 Sparse AutoEncoders
The linear representation hypothesis posits that LLMs linearly represent concepts in neuron activations (Park et al., 2024). However, neuron activations are notoriously superpositioned (Elhage et al., 2022), SAEs enhance the interpretability of LLMs by writing neuron activations as a linear combination of sparse features (Bricken et al., 2023). This reduces the superposition and restores the interpretability (Cunningham et al., 2023). SAEs have recently been applied both in mechanistic interpretability (Nanda et al., 2023; Conmy et al., 2023; Marks et al., 2024) of LLMs, and in deep learning more broadly (Chen and Guo, 2023). SAEs have been scaled to medium size and large Language Models (LMs), such as GPT4 (Templeton et al., 2024; Gao et al., 2024).
@@ 1.2 Our contributions
  • We apply a SAE (EleutherAI, 2024) to Llama 3 8B (Dubey et al., 2024) and extract interpretable features from SEC reports.
@@
  • We use a Logistic Classifier to construct a graph of stocks where similar stocks are connected, and use connected components as clusters with results comparable to or better than standard industry classifications.
  • We demonstrate that a combined model, integrating SAE features with SIC codes, significantly outperforms standalone approaches in explaining pairwise correlations, and hence show that our method can enhance industry classifications.
  • We validate the applicability of our features via cointegration pairs trading strategies, and show that a strategy based on sparse features leads to higher cumulative returns, demonstrating the potential for wider uses within finance.
@@ 2.1 Dataset
Publicly listed companies in the U.S. submit annual financial performance reports to the Securities and Exchange Commission (SEC) which include information on a company’s operations, such as product specifics, subsidiaries, competition, and other financial details(SEC, 2023) . We consider 220,275 annual reports from 1993 to 2020, ignoring any de-lists, accompanied by related meta-data on Company Name, Year, SIC code, and CIK number (a unique SEC corporation identifier). CIK numbers are mapped to their corresponding publicly traded ticker symbol, from which the monthly logged returns are retrieved via Yahoo Finance (2024). Due to a lack of publicly available mappings between CIK codes and GICS classifications, we utilize: (1) SIC codes, and (2) the major group categorization (U.S. Occupational Safety and Health Administration, 2001) , derived from the first two digits of the SIC code, referred to hereafter as BISC (Broader Industry Sector Code). We remove entries with missing or very short company descriptions, ticker information, or monthly returns. Next, company descriptions are tokenized using Meta’s Llama 3 8B Tokenizer (Dubey et al., 2024). We further refine the dataset by retaining only annual reports that have token counts within the context window, and are consistently available for at least five years, leaving us with a dataset of 27,722 reports. In our analysis, we ignore pre-1996 as the sample size is too small. We employ forward testing in our experiments: when any training is done on year $t$ we report and use the predicted results on $t+1$. Our dataset analysis finds that only 0.24% of subsequent company descriptions are identical, with 0.54% within a normalized Levenshtein distance of 0.01, and 12.34% within 0.05. Forward-testing is more robust than out-of-sample testing (Pardo, 1992), and since the descriptions are not repeated across years this supports the generalization of the model.
@@ 2.2 Feature summing
In this work, we face the challenge of comparing sequences of sparse features of arbitrary length (the company descriptions). While this is an emerging topic with limited consensus on best practices, max-pooling has been suggested as a baseline approach for feature aggregation in such settings (Bricken et al., 2024). However, motivated by the specific demands of financial sequence modeling, we propose an alternative approach: sparse feature summing across tokens. This method provides a magnitude-scaled count of how frequently a feature appears within a sequence, reflecting both the number of tokens a feature is active for and its intensity (Lan et al., 2024).
Our approach is inspired by analogous methodologies in literature. For example, the work of Loughran and McDonald highlights the value of summing word counts in financial text analysis to derive domain-relevant insights (Loughran et al., 2009).
We sum sparse features, across tokens, obtained at the layer occurring 90% of the way through the model. At this layer, we capture relevant features from preceding layers via the skip connection(Vaswani et al., 2017), but not the logic-related features that tend to occur at the very last layers (Ghilardi et al., 2024).
The skip connection ensures that a single SAE captures the entire residual stream (Longon, 2024), inherently including information from all preceding layers. This ensures that the summed sparse features represent a comprehensive aggregation of the input sequence’s processing across the model.
We analyze summed sparse features, and observe an interesting exponential decay pattern in feature activation frequencies, which was that approximately half of the features never activate, a fourth activate once, an eighth activate twice, and so forth (Figure 1).
Figure 1 highlights the sparsity of LLM latent features - even when summed across thousands of tokens - that motivates feature summing as an approach. In this context, before summing, a single active feature has on average a value $\approx{0.7}$ (the first bulge).
@@ 2.3 Computing used
2.4 Logistic Classifier
To validate the meaningfulness of the extracted features , we train a logistic regression model aimed at predicting the probability $p_{i,j}$ that two companies $i$ and $j$ are similar, measured by the correlation of $i$ and $j$’s monthly returns [^1] By the Efficient Market Hypothesis, the stock prices of two entirely identical companies move in tandem., opting to use them over higher frequency returns to reduce noise from idiosyncratic events (Huang et al., 2009) [^2] Monthly returns smooth out transient dynamics, providing a clearer view of structural alignment..
For each pair of companies $i$ and $j$, we compute an interaction feature vector $\mathbf{x}_{i,j}$ by performing element-wise multiplication of their summed sparse feature vectors $\mathbf{f}_{i}$ and $\mathbf{f}_{j}$. The resulting interaction vector $\mathbf{x}_{i,j}$ captures the degree of alignment between the features of the two companies, effectively encoding their similarity.
We devise a custom loss function that penalizes the model when the predicted similarity does not align with the actual correlation of their returns. We transform the Pearson correlation coefficients $\rho_{i,j}$ of the companies’ monthly returns to similarity scores $y_{ij}$ in the range $[0,1]$: $y_{i,j}=\frac{\rho_{i,j}+1}{2}$.
Since our features are semantically representative, we expect a strong relationship between the predicted probabilities $p_{i,j}$ and the actual correlations $\rho_{i,j}$ of the companies’ returns. To evaluate this, we train two Logistic models: (1) a model (LC-T1) trained using data from year $t$ to predict year $t+1$, and (2) a model (LC-T3) trained using data from years $t-2$, $t-1$, and $t$ to predict year $t+1$.
Siamese Network (SN): To capture non-linear relationships within the features, we implement a Siamese Neural Network trained using a custom similarity loss function. [^3] Siamese networks are designed to preserve pairwise relationships by mapping inputs into a shared feature space, where the similarity between outputs is maximized for similar pairs and minimized for dissimilar ones.. Each sub-network in the Siamese architecture consists of two fully connected layers with ReLU activation functions: $\text{E}=\text{ReLU}(\text{Lin}(128))\rightarrow\text{ReLU}(\text{Lin}(64))$
After obtaining the 64-dimensional embeddings $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ for a pair of companies, we compute their cosine similarity. Finally, to align predicted similarities with actual financial similarities derived from Pearson correlations $\rho$, we use Mean Squared Error (MSE) loss.
@@ 2.5 Clustering
We test whether our sparse features outperform SIC and BISC through comparison of pairwise correlation of monthly logged returns.
Sparse features lack sufficient locality (unlike embeddings) and smoothness (TopK introduces strong discontinuity) to define reliable similarity metrics. Moreover, half of the features never activate (See Figure 1) and so are unrelated to finance, but are present due to the SAE being trained on a generic dataset (Weber et al., 2024). Thus, we cannot directly cluster the sparse features - we instead employ graph-based clustering methods on thresholds. Unlike traditional clustering, which relies on feature space proximity, our approach constructs undirected graphs $G_{k}$ where nodes represent companies, and edges connect pairs $(i,j)$ that exceed a model-specific threshold [^4] The threshold is optimized using forward testing via temporal cross-validation. Nearby thresholds yield similar results, indicating stability in the clustering performance.. Clusters are the connected components of these graphs, optimized to capture transitive relationships between companies.
@@
(1) $G_{\text{C-T1}}$, where the threshold is the out-of-sample predicted probability $p_{i,j}$ from the Logistic Classifier (LC-T1) trained on year $t$ to predict $t+1$;
(2) $G_{\text{C-T3}}$, where the threshold is the out-of-sample predicted probability $p_{i,j}$ from LC-T3;
(3) $G_{\text{C-TA}}$, based on a second scalar, $A_{i,j}$, derived from PCA-transformed sparse features. The sparse feature vectors $\mathbf{X}$ (131k dimensions) are reduced to 4,000 dimensions using PCA, fitted globally across the years 1996 to 2020 to ensure consistent eigenvectors. For each pair of companies $i$ and $j$, $A_{i,j}$ is computed as:
$A_{i,j}=\sum_{k=1}^{4000}\left|\text{PCA}_{k}(\mathbf{f}_{i})-\text{PCA}_{k}(\mathbf{f}_{j})\right|,$
where $A_{i,j}$ serves as a measure of dissimilarity, with smaller values indicating greater alignment;
(4) $G_{\text{C-TS}}$, uses predicted similarity scores from our Siamese Neural Network;
[…]
(6) $G_{\text{C-LC1\_SIC}}$, derived from predicted similarity scores obtained by running LC-T1 on the one-hot encoded representation of SIC codes; and
(7) $G_{\text{C-LC1\_BISC}}$, similarly derived using the one-hot encoded representation of BISC codes.
We evaluate each clustering model using intra-cluster correlation within $C_{k}$, the group/set-of-clusters for each method. The mean intra-cluster correlation, assessing coherence in financial returns, is defined as:
@@
where $\rho(x,y)$ is the Pearson correlation of logged returns for companies $x$ and $y$. This metric measures the coherence of return behavior within each cluster, reflecting the effectiveness of each clustering method.

### v2

@@ Abstract
Determining company similarity is a vital task in finance, underpinning hedging, risk management, portfolio diversification, and more. Practitioners often rely on sector and industry classifications to gauge similarity, such as SIC-codes and GICS-codes – the former being used by the U.S. Securities and Exchange Commission (SEC), and the latter widely used by the investment community. Since these classifications can lack granularity and often need to be updated, using clusters of embeddings of company descriptions has been proposed as a potential alternative, but the lack of interpretability in token embeddings poses a significant barrier to adoption in high-stakes contexts. Sparse Autoencoders (SAEs) have shown promise in enhancing the interpretability of Large Language Models (LLMs) by decomposing LLM activations into interpretable features. We apply SAEs to company descriptions, obtaining meaningful clusters of equities in the process. We benchmark SAE features against SIC-codes, Industry codes, and Embeddings. Our results demonstrate that SAE features not only replicate but often surpass sector classifications and embeddings in capturing fundamental company characteristics. This is evidenced by their superior performance in correlating monthly returns – a proxy for similarity – and generating higher Sharpe ratio co-integration strategies, which underscores deeper fundamental similarities among companies.
@@
Marco Molinari*, 1and Victor Shao*, 1and Vladimir Tregubiak*, 1 Abhimanyu Pandey1  and Mateusz Mikolajczak1  and Sebastião Kuznetsov Ryder Torres Pereira1 1 LSE.AI, London School of Economics and Political Science https://lseai.org Correspondence: m.molinari1@lse.ac.uk *Equal Contribution
@@ 1 Introduction
Traditional modes of company comparisons rely on relative returns and discrete classifications, e.g. SIC codes or the Global Industry Classification System (GICS), which categorizes companies into 11 sectors and 163 sub-industries (MSCI, 2020). Relying on return spreads can be effective but is not foolproof, as market volatility, economic shifts, business fundamental changes, and temporal factors can disrupt them (Loretan and English, 2000).
Additionally, systems like GICS are limited, as the restricted granularity of a discrete classification system limits dynamic interpretations of companies’ operations, as it falls short in accounting for the duality of certain companies, particularly in emerging industries (Winton, 2018).
This is particularly important for pairs trading, a market-neutral strategy based on mean-reverting return spreads (Ehrman, 2012). Employing a pairs trading strategy with fundamentally similar companies whose returns are co-integrated could lessen the risk of deviation from historical trends (Raghava and Bharadwaj, 2014).
Clustering embeddings of company descriptions has been proposed as a measure of similarity (Vamvourellis et al., 2023), but token embeddings are not interpretable, and this uncertainty is undesirable in the financial sector.
SAEs have the potential to provide an efficient measure of company similarity by decomposing large amounts of financial data into interpretable features (Chen et al., 2020). Moreover, SAEs have recently been applied to LLMs resulting in interpretable decompositions of neural activations (Huben et al., 2024). This motivates their application to textual company descriptions. We are the first, to the best of our knowledge, to compute company similarity using SAEs and SEC filings.
@@ 1.1 Sparse AutoEncoders
The linear representation hypothesis posits that LLMs linearly represent concepts in neuron activations (Park et al., 2024). However, neuron activations are notoriously superpositioned (Elhage et al., 2022), SAEs enhance the interpretability of LLMs by writing neuron activations as a linear combination of sparse features (Bricken et al., 2023). This reduces the superposition and restores the interpretability (Huben et al., 2024). SAEs have recently been applied both in mechanistic interpretability (Nanda et al., 2023; Conmy et al., 2023; Marks et al., 2024) of LLMs, and in deep learning more broadly (Chen and Guo, 2023). SAEs have been scaled to medium size and large Language Models (LMs), such as GPT4 (Templeton et al., 2024; Gao et al., 2024).
@@ 1.2 Our contributions
  • We apply an open source SAE (EleutherAI, 2024) to Llama 3 8B (Dubey et al., 2024), and extract interpretable features from SEC filings.
@@
  • We train a Siamese Network to obtain expected correlations from company descriptions, and construct a minimum spanning tree of stocks where correlated stocks are connected. We then perform hierarchical clustering with results comparable to, or better, than Embeddings, SIC-codes, and Industry categorizations.
  • We demonstrate that a combined model, integrating SAE features with SIC-codes and Industry categorizations, significantly outperforms standalone approaches in explaining pairwise correlations, showing that our method can enhance industry classifications.
  • We validate the applicability of our features via co-integration pairs trading strategies, and show that a strategy based on sparse features leads to a higher Sharpe ratio, demonstrating the potential for wider uses within finance.
@@ 2.1 Dataset
Publicly listed companies in the U.S. submit annual financial performance reports to the Securities and Exchange Commission (SEC) which include information on a company’s operations, such as product specifics, subsidiaries, competition, and other financial details(SEC, 2023) . We consider 220,275 annual reports from 1993 to 2020, ignoring any de-lists, accompanied by related meta-data on Company Name, Year, SIC-code, and CIK number (a unique SEC corporation identifier). CIK numbers are mapped to their corresponding publicly traded ticker symbol, from which the monthly logged returns are retrieved via Yahoo Finance (2024). Due to a lack of publicly available mappings between CIK codes and GICS classifications, we utilize: (1) SIC-codes, and (2) the industry/major-division categorization [^1] The first 3 digits of the SIC code splits companies into 12 industry/major-divisions, referred to hereafter as BISC (Broader Industry Sector Code). (U.S. Occupational Safety and Health Administration, 2001). We remove entries with missing or very short: company descriptions, ticker information, or monthly returns. Next, company descriptions are tokenized using Meta’s Llama 3 8B Tokenizer (Dubey et al., 2024). We further refine the dataset by retaining only annual reports that have token counts within the context window, and are consistently available for at least five years, leaving us with a dataset of 27,722 reports. In our analysis, we ignore pre-1996 as the sample size is too small.
We employ forward testing in our experiments: when any training is done on year $t$, we report and use the predicted results on $t+1$. Our dataset analysis finds that only 0.24% of subsequent company descriptions are identical, with 0.54% within a normalized Levenshtein distance of 0.01, and 12.34% within 0.05. Forward-testing is more robust than out-of-sample testing (Pardo, 1992), and since the descriptions are not repeated across years, the model remains generalized.
@@ 2.2 Feature summing
In this work, we face the challenge of comparing sparse feature sequences of arbitrary lengths, where best practices are not well-established, though max-pooling has been proposed as a baseline for feature aggregation (Bricken et al., 2024). However, motivated by the specific demands of financial sequence modeling, we propose an alternative, employing sparse feature summing across tokens. This method provides a magnitude-scaled count of how frequently a feature appears within a sequence, reflecting both the number of tokens a feature is active for and its intensity (Lan et al., 2024).
Our approach is inspired by analogous methodologies in literature. For example, Loughran et al. (2009) highlight the value of summing word counts in financial text analysis to derive domain insights.
We sum sparse features, across tokens, from an SAE (EleutherAI, 2024) applied at the layer occurring 90% of the way through the model. At this layer, we capture relevant features from preceding layers via the skip connection (Vaswani et al., 2017), but not the logic-related features that tend to occur at the very last layers (Ghilardi et al., 2024).
The skip connection ensures that a single SAE captures the entire residual stream (Longon, 2024), inherently including information from all preceding layers, thus ensuring that the summed sparse features represent a comprehensive aggregation of the descriptions’s representation in the model.
We analyze summed sparse features, and observe an interesting exponential decay pattern in feature activation frequencies, where half the features never activate, a fourth activate once, an eighth activate twice, and so forth (Figure 1).
Figure 1 highlights the sparsity of LLM latent features – even when these are summed across thousands of tokens – motivating feature summing as an approach. In this context, before summing, a single active feature has, on average, a value of $\approx{0.7}$ (the first bulge).
@@
2.3 Models
2.3.1 Embedders
As a baseline, we replicate Vamvourellis et al. (2023)’s embedding methodology and obtain embeddings for company descriptions in our dataset. In particular, we use:
  1. 1.
    BERT: Given the maximum token length of 512, we employ chunking by splitting company descriptions into overlapping chunks of 512 tokens, extracting CLS embeddings for each chunk, and averaging these embeddings to generate a single 1536-token representation, following (Vamvourellis et al., 2023).
  2. 2.
    Sentence-BERT (SBERT): Building on BERT, SBERT improves latency substantially (Reimers and Gurevych, 2019) and encodes meaning on the more abstract sentence level. We again employ chunking to 1536 tokens.
  3. 3.
    PaLM-gecko: Pathways Language Model (PaLM) (Chowdhery et al., 2022).
2.4 Siamese Networks
Given the comparative nature of our use case, we employ Siamese neural networks (Bromley et al., 1993) to validate the meaningfulness of the extracted features. Specifically, the Siamese network is trained to predict the correlation $p_{i,j}$ between two companies $i$ and $j$  [^2] Monthly returns are used over higher-frequency returns to reduce noise from idiosyncratic events (Huang et al., 2009)..
The Siamese network uses similarity information to position correlated companies closer together in the feature space. It employs two identical sub-networks with shared weights and parameters (Koch et al., 2015). Features extracted by the SAE from companies $A$ and $B$, denoted as $f_{1}$ and $f_{2}$, are passed through two dense layers, each followed by a ReLU activation, to produce encoded representations $z_{1}$ and $z_{2}$. Cosine similarity of $z_{1}$ and $z_{2}$ is then used in the MSE loss function for network optimization [^3] See Appendix E. We train a new Siamese network for each year, using data from $t$ (1993–2019) to predict $t+1$ (1994–2020), iterating year by year.
@@ 2.5 Clustering
We benchmark our sparse features against embeddings and SIC/BISC-codes, where each SIC/BISC-code is its own cluster.
@@
where $\rho(x,y)$ is the Pearson correlation of logged monthly returns for companies $x$ and $y$. This metric quantifies stock return coherence within clusters, evaluating each clustering method’s effectiveness.
2.5.1 Clustering Embeddings
Each of the embedders we employ is used to define a group/set-of-clusters ($G_{\text{embedder}}$): (a) $G_{\text{BERT}}$; (b) $G_{\text{SBERT}}$; (c) $G_{\text{PaLM-gecko}}$ following Vamvourellis et al. (2023). We first reduce the dimensionality of the embeddings using UMAP (McInnes et al., 2020), then generate clusters through Spectral Clustering. See Appendix A for methodology details.
2.5.2 Clustering Sparse Features
Sparse features lack the locality and smoothness of embeddings (Kiros et al., 2015; Bischke et al., 2019) to define reliable similarity metrics. Moreover, half of the features never activate (See Figure 1) and hence are unrelated to finance, but are present due to the SAE being trained on a generic dataset (Weber et al., 2024). Thus, we cannot directly cluster the sparse features.
To cluster the sparse features, we adopt the graph-theoretic framework of Bonanno et al. (2004), employing Minimum Spanning Trees (MSTs) to extract hierarchical structures from financial data. A fully connected graph is constructed with edge weights representing a particular distance metric (see metrics 1, 2 below). The MST encodes a subdominant ultrametric, with ultrametric distance defined by the maximum edge weight on the unique path between two nodes [^4] To enforce the ultrametric property, we employ single-linkage hierarchical clustering, which groups nodes by iteratively merging the pair of clusters with the smallest maximum distance between any two points. This process inherently satisfies the ultrametric inequality $(d_{ij}\leq\max(d_{ik},d_{kj}))$ by construction.. We remove edges above a specified weight level [^5] See Appendix B on edge-weight cutoff optimisation, generating clusters directly from the MST. This eliminates the need for additional clustering steps, ensuring stable and interpretable results consistent with Bonanno et al. (2004).
Metric 1: Siamese-derived Metric: This cluster group is constructed using a siamese-derived ultrametric, and will be called $G_{\text{C-TS}}$. First we obtain predicted correlations from the Siamese (SN) model. Then, following (Bonanno et al., 2004) we use these to define a distance metric between stock pairs, where $\Delta t$ is the time horizon:
$d_{i,j}(\Delta t)=\sqrt{2(1-\rho_{i,j}(\Delta t))}.$
The distance matrix $\mathbf{D}(\Delta t)$ [^6] With this choice, $d_{i,j}(\Delta t)$ fulfills the three axioms of a metric: (i) $d_{i,j}(\Delta t)=0$ iff $i=j$; (ii) $d_{i,j}(\Delta t)=d_{j,i}(\Delta t)$; (iii) $d_{i,j}(\Delta t)\leq d_{i,k}(\Delta t)+d_{k,j}(\Delta t)$. is then used to determine the MST connecting the $n$ stocks.
Metric 2: Manhattan Metric: This group, $G_{\text{C-TM}}$, is based on a second metric, $M_{i,j}$, derived from PCA-transformed sparse features, where 131k-dimensional vectors are reduced to 4,000 dimensions [^7] We fit PCA globally across 1996–2020 for consistent eigenvectors.. For each pair of companies $i$ and $j$, $M_{i,j}$ is computed as:
$M_{i,j}=\sum_{k=1}^{4000}\left|\text{PCA}_{k}(\mathbf{f}_{i})-\text{PCA}_{k}(\mathbf{f}_{j})\right|$
where $M_{i,j}$ serves as a metric888 See Appendix C for the full proof of dissimilarity (Kotsiantis et al., 2001).